OpenStudy (anonymous):

A juggler has a bag containing three yellow balls, five green balls, three blue balls, and one red ball, all the same size. The juggler pulls a ball from the bag at random. Then, without replacing it, he pulls out a second ball. What is the probability that the juggler first draws a green ball followed by a blue ball?

OpenStudy (anonymous):

8/12 or 2/3

OpenStudy (kirbykirby):

Initially there are (3+5+3+1)=12 balls total. So, the probability of getting a green ball is 5/12. Now 1 ball has been removed, so there are only 11 left. The # of blue balls is the same, so the probability of drawing a blue ball is 3/11 So the probability of getting 1 green ball AND 1 blue ball after is (you must use the product rule): \[ \frac{5}{12}\cdot\frac{3}{11}=\frac{15}{132}=\frac{5}{44}\]

OpenStudy (anonymous):

o-o